Evaluate Gamma (-frac{3}{2})
1
Expert's answer
2019-09-16T10:55:59-0400

"\Gamma(-\frac{3}{2})=\int\limits_0^\infty x^{-\frac{3}{2}-1}e^{-x}dx"

It is known that

"\Gamma(1-z)\Gamma(z)=\frac{\pi}{sin\pi z}"

Hence for z=1/2

"\Gamma^2(\frac{1}{2})=\pi"

"\Gamma(\frac{1}{2})=\sqrt{\pi}"

It is also known that "\Gamma(z+1)=z\Gamma(z)"

Therefore

"\Gamma(\frac{1}{2})=-\frac{1}{2}\Gamma(-\frac{1}{2})=\frac{3}{4}\Gamma(-\frac{3}{2})"

"\Gamma(-\frac{3}{2})=\frac{4}{3}\Gamma(\frac{1}{2})=\frac{4\sqrt{\pi}}{3}"